# Searching for coupled, hyperlight scalars across cosmic history

### v1

@@ B.2 Impact of BBN consistency
In the class of models we consider (Section II), shifts in the fundamental constants during BBN are generally the same as during recombination (unless the matter effects of Section II.3 are important, in which case it is most likely that the constants differ from their present-day values even more at earlier times). Following the discussion in Ref. [60], the primary effects of the helium yield on the CMB anisotropies are on diffusion damping rate and the width of the visibility function. The change to the helium yield predicted by BBN in particular introduces implicit dependence on $m_{e,i}$ per Eq. 90 that (in principle) breaks the degeneracy of $m_{e,i}$ with $\Lambda$CDM parameters.
@@
Figure 10: Marginalized posterior distributions over the early-time electron mass (left) and fine-structure constant (right), each varied independently, for the corresponding scalar field models with (red) and without (blue) accounting for the effect of the variation in fundamental constants on BBN. The posteriors only use the Planck 2018 likelihoods, since the difference is negligible when including data that better constrain the models. Vertical lines mark the median and $16$th and $84$th percentiles; the median and corresponding ±$\pm 1\sigma$ uncertainties for each parameter are reported above each panel.

### v2

@@ B.2 Impact of BBN consistency
In the class of models we consider (Section II), shifts in the fundamental constants during BBN are generally the same as during recombination (unless the matter effects of Section II.3 are important, in which case for positive couplings the constants differ from their present-day values even more at earlier times). Following the discussion in Ref. [60], the primary effects of the helium yield on the CMB anisotropies are on diffusion damping rate and the width of the visibility function. The change to the helium yield predicted by BBN in particular introduces implicit dependence on $m_{e,i}$ per Eq. 90 that (in principle) breaks the degeneracy of $m_{e,i}$ with $\Lambda$CDM parameters.
@@
Figure 10: Marginalized posterior distributions over the early-time electron mass (left) and fine-structure constant (right), each varied independently, for the corresponding scalar field models with (red) and without (blue) the effect of the variation in fundamental constants on BBN accounted for. The posteriors only use the Planck 2018 likelihoods, since the difference is negligible when including data that better constrain the models. Vertical lines mark the median and 161616th and 848484th percentiles; the median and corresponding ±$\pm 1\sigma$ uncertainties for each parameter are reported above each panel.